OpenStudy (anonymous):

Its TIME To have a ERICeaglemathsriddle Solve it: \[I=\int\limits_{0}^{a}f(x)dx\] \[a>0\] \[and, f(x)<0 \] for all \[0\le x \le a\]

OpenStudy (lgbasallote):

lol do you even know what you're asking?

OpenStudy (anonymous):

ya I know??

OpenStudy (anonymous):

@lgbasallote @ramkrishna @ParthKohli @chand @dumbcow pls help with it

OpenStudy (dumbcow):

well the answer will be negative since f(x) <0 for all x but can't integrate a function i know nothing about

OpenStudy (anonymous):

come on @dumbcow help

OpenStudy (dumbcow):

thats all i got...you stumped me , please tell me what the solution is

OpenStudy (anonymous):

I'm asking to you my bro!

OpenStudy (anonymous):

@nbouscal

OpenStudy (dumbcow):

if you are looking for numerical answer...there is none you can use fundamental thm of calculus \[\int\limits_{0}^{a} f(x) dx = F(a) - F(0)\] where F'(x) = f(x)

OpenStudy (anonymous):

since F(a) and F(0) both are less than zero so if |F(a)|<|F(0)| then I > 0 and vice versa.